What is the instantaneous slope of y = eight over x at x = –3?
negative 8 over 3
8 over 3
negative 8 over 9
8 over 9

Answer by Cromlix(4381)

Hi there,
The gradient at x = -3 on
the equation y = 8/x
y = 8/x
y = 8x^-1
dy/dx = -8x^-2
dy/dx = -8/x^2
@ x = -3
= -8/(-3)^2
= -8/9
Hope this helps :-)